Circular No. 7779

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 11D (TEMPEL-SWIFT) = P/2001 X3 (LINEAR) C. Hergenrother, Lunar and Planetary Laboratory; and K. Muraoka,

Kochi, Japan, suggested a link between comet 11D (last seen in 1908) and P/2001 X3 (cf. IAUC 7778) — a linkage confirmed at the Minor Planet Center and by S. Nakano (Sumoto, Japan). The orbital elements below are by Nakano (from 43 observations, 1908–2001, mean residual 0′′.8; nongravitational parameters A_{1} = +0.13 ± 0.01, A_{2} = −0.0134 ± 0.0007). The comet was not found in 1963 despite a prediction by B. G. Marsden (IAUC 1838, 1839, 1840). More recent predictions were made by Marsden and Sekanina (1971, A.J. 76, 1142), by Nakano (Comet Handbooks for 1989, 1995, and 1996, Oriental Astronomical Association; and NK 686), and by Muraoka (Comet Handbook for 2001, OAA). The indicated correction to Nakano’s 2001 prediction (1998, NK 686) is ΔT = +3.4 days.

Epoch = 2002 Jan. 6.0 TT

T = 2001 Dec. 30.75519 TT  ω = 163.^{o}63401
e = 0.5391227  Ω = 240.71112 }2000.0
q = 1.5840469 AU  i = 13.46049

a = 3.4370255 AU n^{o} = 0.15467850 P = 6.372 years

Epoch = 1908 Sept. 16.0 TT

T = 1908 Oct. 5.07032 TT  ω = 113.^{o}40868
e = 0.6381330  Ω = 291.84116 }2000.0
q = 1.1524985 AU  i = 5.44909

a = 3.1848677 AU n^{o} = 0.17340714 P = 5.684 years

SUPERNOVA 2001if NEAR MCG +06-6-43 S. Benetti, G. Altavilla, A. Pastorello, M. Turatto, S. Desidera, E.

Giro, C. Pernechele and A. Bortolussi, Osservatorio Astronomico di Padova; and E. Cappellaro, Osservatorio Astronomico di Capodimonte, write: “A spectrum (range 400–780 nm, resolution 2.5 nm) of SN 2001if (cf. IAUC 7771), obtained on Dec. 19.87 UT with the Asiago 1.82-m telescope (+ AFOSC), shows it to be a type-Ia supernova 15 ± 5 days after maximum light. The SN spectral features and age are consistent with those of a type-Ia supernova having a photospheric expansion velocity of ∼ 9000 km/s (deduced from the Si II 635.5-nm feauture) at a recession velocity of ∼ 11 400 km/s. Thus, SN 2001if belongs to the MCG +06-6-43 cluster, even if the SN is not clearly associated with any galaxy of the field.”

2001 December 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green